Ok here is the equation 2(x-4)-(x+2)=(x+2)(x-4)
I am supposed to solve for x since none of the other equations work I am to put it into ax^2+bx+c=0 form. My question is how does it go from 2(x-4)-(x+2)=(x+2)(x-4) to x^2-3x+2=0? I know the rest of the problem this is the only part I am having difficulty understanding.

Answer by richwmiller(17219)

2(x-4)-(x+2)=(x+2)(x-4)
2x-8-x-2=x^2-2x-8
x-10 = x^2-2x-8
0=x^2-3x+2